OpenStudy (anonymous):

-7x+4=-1x-8+5 (2-3x)

OpenStudy (anonymous):

x=−2/9

OpenStudy (anonymous):

-7x+4=-1x-8+5 (2-3x) -7x+4=-x-8+10-15x -7x-x+15x=-8+10-4 9x/9=-2/9 ans -2/9